Explain rotational energy in term Of J.


1
Expert's answer
2021-01-27T09:31:35-0500

In molecular physics and thermodynamics, one degree of freedom of rotational motion has energy "\frac{1}{2}kT" . At room temperature "(298\ K)"


"\frac{1}{2}kT\approx\frac{1}{2}\cdot 1.38\cdot10^{-23}\cdot298=206\cdot10^{-23}\ (\ J\ )" .Answer